### v1

# Constraints on the Hubble and matter density parameters with and without modelling the CMB anisotropies

Abstract
We consider constraints on the Hubble parameter $H_{0}$ and the matter density parameter $\Omega_{\mathrm{M}}$ from: (i) the age of the Universe based on old stars and stellar populations in the Galactic disc and halo (Cimatti & Moresco, 2023); (ii) the turnover scale in the matter power spectrum, which tells us the cosmological horizon at the epoch of matter-radiation equality (Philcox et al., 2022); and (iii) the shape of the expansion history from supernovae (SNe) and baryon acoustic oscillations (BAOs) with no absolute calibration of either, a technique known as uncalibrated cosmic standards (UCS; Lin, Chen, & Mack, 2021). A narrow region is consistent with all three constraints just outside their $1\sigma$ uncertainties. Although this region is defined by techniques unrelated to the physics of recombination and the sound horizon then, the standard Planck fit to the CMB anisotropies falls precisely in this region. This concordance argues against early-time explanations for the anomalously high local estimate of $H_{0}$ (the ‘Hubble tension’), which can only be reconciled with the age constraint at an implausibly low $\Omega_{\mathrm{M}}$. We suggest instead that outflow from the local KBC supervoid (Keenan, Barger, & Cowie, 2013) inflates redshifts in the nearby universe and thus the apparent local $H_{0}$. Given the difficulties with solutions in the early universe, we argue that the most promising alternative to a local void is a modification to the expansion history at late times, perhaps due to a changing dark energy density.
Cosmology (343) – Hubble constant (758) – Matter density (1014) – Cosmological parameters from large-scale structure (340) – Stellar ages (1581)
journal: The Open Journal of Astrophysics
[…]
1 Introduction
The Universe is thought to have undergone an early period of rapid accelerated expansion known as inflation, which would cause any initial departure from flatness to rapidly decay (Guth, 1981). After this inflationary era and the subsequent era of radiation domination, a flat FLRW universe (Friedmann, 1922, 1924; Lemaître, 1931; Robertson, 1935, 1936a, 1936b; Walker, 1937) is largely governed by just two parameters: the fraction of the cosmic critical density in the matter component ($\Omega_{\mathrm{M}}$) and $h$, the Hubble constant $H_{0}$ in units of 100 km/s/Mpc. It is well known that if we assume the standard cosmological paradigm known as Lambda-Cold Dark Matter ($\Lambda$CDM; Efstathiou, Sutherland, & Maddox, 1990; Ostriker & Steinhardt, 1995), these parameters can be obtained to high precision from the anisotropies in the cosmic microwave background (CMB; e.g. Planck Collaboration VI, 2020). However, the inferred $h$ is inconsistent with the locally inferred value from the redshift gradient $z^{\prime}\equiv dz/dr$, which measures how quickly the redshift $z$ rises with distance $r$. In a homogeneously expanding universe, the local $z^{\prime}=H_{0}/c$, where $c$ is the speed of light. The anomalously high local $cz^{\prime}$ is known as the Hubble tension (e.g. Di Valentino, 2021; Perivolaropoulos, 2024).
It has been argued that $h$ and $\Omega_{\mathrm{M}}$ are some of the most important cosmological parameters to consider when trying to address the Hubble tension (Toda et al., 2024). Some proposed solutions modify the physics prior to recombination at $z=1100$ but have minimal impact beyond the first Myr of cosmic history. This motivates us to infer $h$ and $\Omega_{\mathrm{M}}$ without modelling the CMB anisotropies in $\Lambda$CDM -- such results may not be applicable to alternative cosmological models, especially those involving early dark energy (EDE; e.g. Poulin et al., 2019). We therefore focus on three non-CMB constraints, two of which are almost completely immune to the physics in the first Myr. The third assumes $\Lambda$CDM prior to and shortly after matter-radiation equality at $z=3400$. This is because early-time solutions to the Hubble tension often modify the physics only shortly prior to recombination, thus not much affecting the radiation-dominated era when the universe was 7$\gtrsim$ 7$\times$ 7 × younger (see Appendix A). All three constraints should therefore remain valid in such alternative cosmologies.
In Section 2, we explain the techniques used to constrain $H_{0}$ and $\Omega_{\mathrm{M}}$. We then present our results and discuss them in Section 3 before concluding in Section 4.
2 Methods
In what follows, we briefly describe each of the three considered non-CMB constraints on $\Omega_{\mathrm{M}}$ and $H_{0}$. For completeness, we also consider the standard Planck constraint and the local $cz^{\prime}$.
2.1 Uncalibrated cosmic standards (UCS)
There is much controversy about the absolute magnitude of SNe and the comoving size $r_{d}$ of the BAO ruler. Leaving these as free parameters in the UCS technique, we can still constrain the shape of the expansion history and thus $\Omega_{\mathrm{M}}$ (Lin, Chen, & Mack, 2021). [^1] Those authors also use the angular scale of the first peak in the CMB power spectrum as an extra angular BAO data point, but this does not assume any pre-recombination model. The size of the CMB ruler is $r_{\star}=r_{d}/1.0184$, a ratio which is essentially the same even in quite non-standard cosmological models (see section 2.2 of Vagnozzi, 2023). This yields the tight constraint $\Omega_{\mathrm{M}}=0.302\pm 0.008$.
2.2 The horizon scale at matter-radiation equality
The cosmic expansion history has largely been a power-law due to matter being by far the dominant component, apart from the relatively recent phenomenon of dark energy dominance. The combination of a power-law primordial power spectrum (Harrison, 1970; Zeldovich, 1972) with a power-law expansion history leads to a power-law spectrum also at late times. This behaviour breaks down on small scales because at the early times when these modes entered the cosmic horizon, the universe was still dominated by radiation. During that era, even sub-horizon matter density perturbations could not grow due to the Meszaros effect (Meszaros, 1974). This leads to a characteristic peak in the matter power spectrum at a comoving wavenumber of $k_{\mathrm{eq}}$.
Surveys of large-scale structure in redshift space provide the tight constraint $k_{\mathrm{eq}}/h=\left(1.64\pm 0.05\right)\times 10^{-2}$/Mpc (Philcox et al., 2022). This is important to our discussion because $k_{\mathrm{eq}}/h\propto\Omega_{\mathrm{M}}h$ (equation 3 of Eisenstein & Hu, 1998, note they use $\Omega_{0}$ to denote $\Omega_{\mathrm{M}}$). We obtain the proportionality constant by noting that Philcox et al. (2022) report most likely values of $h=0.648$ and $\Omega_{\mathrm{M}}=0.338$ (see their figure 2). This combination presumably corresponds to their most likely $k_{\mathrm{eq}}/h$. We assume its fractional uncertainty is the same as that in $\Omega_{\mathrm{M}}h$.
2.3 Stellar ages in the Galactic halo
Cimatti & Moresco (2023) compile reliable measurements of the ages of stars, globular clusters, and ultrafaint dwarf galaxies (UFDs) in the Galactic disc and halo. The inverse variance weighted mean age of their 11 considered objects is
$t_{\mathrm{age}}=14.05\pm 0.25\,\text{Gyr}.$ (1)
We assume these objects formed in $t_{\mathrm{f}}=0.2$ Gyr, with $t_{\mathrm{f}}$ uncertain by a factor of 2. The resulting range of $100-400$ Myr corresponds to their estimated range of $11-30$ for the formation redshift. The total age of the Universe is thus
$t_{\mathrm{U}}~{}=~{}t_{\mathrm{age}}+t_{\mathrm{f}}.$ (2)
We obtain $H_{0}t_{\mathrm{U}}$ for each $\Omega_{\mathrm{M}}$ by inverting equation 45 of Haslbauer, Banik, & Kroupa (2020). We note that $H_{0}t_{\mathrm{U}}\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr\propto\cr\kern 2.0pt\cr\sim\cr\kern-2.0pt\cr}}}\Omega_{\mathrm{M}}^{-0.28}$ (equation 35 of Poulin, Smith, & Karwal, 2023).
2.4 The Planck CMB constraint in $\Lambda$CDM
We find the size and orientation of the Planck error ellipse from the published uncertainties on $\Omega_{\mathrm{M}}$ and $h$ (see the TTTEEE column of table 5 in Tristram et al., 2024). The tightest constrained combination is $\Omega_{\mathrm{M}}h^{3}$ (Kable, Addison, & Bennett, 2019). Working in the space of $\left(\ln\Omega_{\mathrm{M}},\ln h\right)$, the short axis of the error ellipse is thus towards the direction $\left(1,3\right)/\sqrt{10}$, with uncertainty $\sigma_{+}$ (named for its positive slope). The long axis is the orthogonal direction $\left(-3,1\right)/\sqrt{10}$, with uncertainty $\sigma_{-}$. Decomposing the $\ln\Omega_{\mathrm{M}}$ and $\ln h$ directions into combinations of these two statistically independent directions, we find $\sigma_{+}$ and $\sigma_{-}$ by solving
$\begin{bmatrix}0.1&~{}0.9\\ 0.9&~{}0.1\end{bmatrix}\begin{bmatrix}\sigma_{+}^{2}\\ \sigma_{-}^{2}\end{bmatrix}~{}=~{}\begin{bmatrix}\sigma^{2}\left(\ln\Omega_{\mathrm{M}}\right)\\ \sigma^{2}\left(\ln h\right)\end{bmatrix},$ (3)
where $\sigma\left(\ln\Omega_{\mathrm{M}}\right)$ is the fractional uncertainty in $\Omega_{\mathrm{M}}$ (similarly for $h$). We then scale the ellipse by a factor $\chi$ so that the enclosed probability is 68.27%, corresponding to the usual $1\sigma$ definition. Since there are two degrees of freedom and we assume Gaussian errors, we find that $\chi=1.52$ by inverting equation 24 of Asencio, Banik, & Kroupa (2021). Thus, the error ellipse in $\ln$-space has semi-minor axis $\chi\sigma_{+}=0.0024$ and semi-major axis $\chi\sigma_{-}=0.036$. The axis ratio of 15 indicates negligible uncertainty on $\Omega_{\mathrm{M}}h^{3}$ compared to other combinations.
2.5 The local redshift gradient
As a representative of the high local estimates of $cz^{\prime}$ (usually equated with $H_{0}$), we show the SH0ES result (Riess et al., 2022). Their study calibrates the lowest rung of the distance ladder using the parallaxes of star clusters rather than individual stars, which are less accurate. We emphasize that many other measurements of the local $cz^{\prime}$ by different teams using different techniques and instruments give quite similar values (e.g., see figure 10 of Riess & Breuval, 2024, and references therein). Indeed, distances obtained by the James Webb Space Telescope (JWST) are in good agreement with those obtained previously by the Hubble Space Telescope (HST; Freedman et al., 2024; Riess et al., 2024). The fundamental plane (FP) relation of elliptical galaxies has recently been used to anchor the local $cz^{\prime}$ to $d_{\mathrm{Coma}}$, the distance to the Coma Cluster (Said et al., 2024). Measurements of $d_{\mathrm{Coma}}$ over the last three decades using a variety of techniques give consistent results, which combined with the above FP result gives a high local $cz^{\prime}$ consistent with other studies (Scolnic et al., 2024). $cz^{\prime}$ can be reduced to the Planck $\Lambda$CDM value only if we assume an implausibly large $d_{\mathrm{Coma}}\gtrsim 110$ Mpc, but observations show that $d_{\mathrm{Coma}}\lesssim 100$ Mpc. This makes it increasingly unlikely that the local $cz^{\prime}$ will one day be revised downwards by the necessary 10%.
3 Results and Discussion
Figure 1: The $1\sigma$ constraints on $\Omega_{\mathrm{M}}$ and $h\equiv H_{0}$ in units of 100 km/s/Mpc from the shape of the expansion history traced by SNe and BAOs with no absolute calibration of either (red; Lin, Chen, & Mack, 2021), the turnover scale in the matter power spectrum (green; Philcox et al., 2022), and the ages of old stars in the Galactic disc and halo (blue; Cimatti & Moresco, 2023), with the light blue band allowing a factor of 2 uncertainty in their formation time. The grey error ellipse shows the Planck fit to the CMB anisotropies in $\Lambda$CDM (Tristram et al., 2024), which provide the tightest constraint on the combination $\Omega_{\mathrm{M}}h^{3}$ (Kable, Addison, & Bennett, 2019). The white dot at its centre shows the most likely values. The yellow band shows $h$ estimated from the local redshift gradient by the SH0ES team (Riess et al., 2022).
Our results are shown in Figure 1. The $1\sigma$ constraints from UCS, $k_{\mathrm{eq}}/h$, and stellar ages are not mutually compatible. These bands enclose a narrow triangle in parameter space that is consistent with all three constraints just outside their $1\sigma$ regions. Remarkably, this is precisely where we find the tight Planck constraint from the CMB anisotropies assuming $\Lambda$CDM. The overlap of these two very small regions of parameter space strongly suggests that the underlying assumptions are largely correct and that the model parameters $\Omega_{\mathrm{M}}$ and $H_{0}$ are in this narrow range.
This concordance highlights the discrepant nature of the local $cz^{\prime}$, which if equated with $H_{0}$ can only be reconciled with the observed age of the Universe at an unrealistically low $\Omega_{\mathrm{M}}$. We suggest instead that $cz^{\prime}$ has been inflated by outflow from the local KBC supervoid (Keenan, Barger, & Cowie, 2013), which is evident across the whole electromagnetic spectrum (see section 1.1 of Haslbauer, Banik, & Kroupa, 2020, and references therein). The model developed by those authors solves the Hubble tension consistently with the observed void density profile. It also successfully predicted the bulk flow curve out to $250/h$ Mpc (Watkins et al., 2023; Mazurenko et al., 2024). The model requires faster structure growth than expected in $\Lambda$CDM, as otherwise local structure cannot solve the Hubble tension through cosmic variance in $H_{0}$ (expected to be only 0.9 km/s/Mpc; Camarena & Marra, 2018). Independently of the proposed local void model, the fact that bulk flows on a scale of $250/h$ Mpc are 4$\approx$ 4$\times$ 4 × the $\Lambda$CDM expectation (Watkins et al., 2023; Whitford et al., 2023) implies that peculiar velocities are larger than expected by about this factor on the relevant scale. This in turn would raise the expected cosmic variance in $H_{0}$ to perhaps 3 km/s/Mpc, which is sufficient to solve the Hubble tension as a $2\sigma$ local fluctuation.
Alternatively, a modification to the expansion history at late times may have little impact on $t_{\mathrm{U}}$ and the angular diameter distance to the CMB. If the modification is carefully tuned, it could enhance the present expansion rate enough to solve the Hubble tension while not much affecting the other observables. Such a modification may arise from dark energy not being a pure cosmological constant (e.g., Harko, 2023; Yao & Meng, 2023). The main problem with this is that a larger $H_{0}$ requires a higher energy density according to the Friedmann equations, which would mean that the dark energy density rises with time (Dahmani et al., 2023). This leads to a phantom equation of state, which suffers from theoretical issues with vacuum stability (Ludwick, 2017). It would also be unusual if we happened to be living at just the time when dark energy starts deviating from a cosmological constant. Moreover, we would have to neglect the evidence that structure formation is more efficient than expected in $\Lambda$CDM, e.g. from the high mass, redshift, and collision velocity of the El Gordo interacting galaxy clusters (Asencio et al., 2021, 2023). A faster background expansion rate would create more Hubble drag on growing structures, making it more difficult to explain the fast observed bulk flows and form large and deep voids. Even so, a modification to the expansion history at late times would avoid the serious difficulties pointed out in this contribution with modifications at early times, even if such modifications can achieve a good fit to the CMB anisotropies (see also Vagnozzi, 2023, and references therein).
4 Conclusions
We consider constraints on $H_{0}$ and $\Omega_{\mathrm{M}}$ from the shape of the cosmic expansion history as traced by SNe and BAO measurements without any absolute calibration of either (UCS; Lin, Chen, & Mack, 2021), the horizon size at matter-radiation equality as deduced from the turnover scale in the matter power spectrum (Philcox et al., 2022), and the age of the Universe from old stars in the Galactic disc and halo (Cimatti & Moresco, 2023). These constraints assume a flat FLRW cosmology but are insensitive to the physics shortly prior to recombination, with the age and UCS constraints being almost completely immune to what happened in the first Myr. We find that the $1\sigma$ regions allowed by these constraints do not overlap, but a narrow region of parameter space is consistent just outside the $1\sigma$ regions defined by these constraints (Figure 1). Remarkably, the standard Planck fit to the CMB anisotropies in $\Lambda$CDM falls precisely in this region (Tristram et al., 2024). This strongly suggests that the anomalously high local $cz^{\prime}$ (the Hubble tension) is not caused by a modification to $\Lambda$CDM at early times in cosmic history. The required late-time modification can either be at the background level through a slight tweak to the expansion history, or it could be due to a local inhomogeneity. Since cosmic variance in the local $cz^{\prime}$ is only 0.9 km/s/Mpc in $\Lambda$CDM (Camarena & Marra, 2018), solving the Hubble tension this way would require enhanced structure formation on 100 Mpc scales, which might then permit the formation of a local supervoid that solves the Hubble tension. There is strong evidence for just such a void (Keenan, Barger, & Cowie, 2013; Haslbauer, Banik, & Kroupa, 2020) and for bulk flows being about 4×4$\times$4 × faster than expected in $\Lambda$CDM on scales relevant to the measurement of the local $cz^{\prime}$ (Watkins et al., 2023; Whitford et al., 2023), which is to be expected in the local void scenario (Mazurenko et al., 2024). Our work adds to the growing evidence that modifications to $\Lambda$CDM at early times in cosmic history are unlikely to solve the Hubble tension, even if it is possible to preserve a good fit to the CMB anisotropies (Vagnozzi, 2023; Toda et al., 2024).
Acknowledgements
IB is supported by Royal Society University Research Fellowship 211046. NS is supported by the AKTION grant number MPC-2023-07755 and the Charles University Grants Agency (GAUK) grant number 94224. The authors thank Phillip Helbig and Harry Desmond for helpful comments.
Appendix A The expansion history around matter-radiation equality
For the first several Gyr of the universe, the expansion history was almost entirely governed by the densities of matter and radiation. The Friedmann equation was thus of the form
$\frac{\dot{a}}{a}~{}=~{}H_{0}\sqrt{\Omega_{\mathrm{M}}\,a^{-3}+\Omega_{\mathrm{r}}\,a^{-4}}\,,$ (A1)
where $a$ is the cosmic scale factor with value 1 today, an overdot denotes a time derivative, and $\Omega_{\mathrm{M}}$ ($\Omega_{\mathrm{r}}$) is the present fraction of the cosmic critical density in matter (radiation). We can solve this by separating the variables and using the substitution $a=a_{\mathrm{eq}}/\tan^{2}\theta$. Applying the usual boundary condition that $a=0$ when $t=0$ gives the following result:
$H_{0}\sqrt{\Omega_{\mathrm{M}}}t~{}=~{}\frac{2\,a_{\mathrm{eq}}^{3/2}}{3}\left[2+\sqrt{\frac{a}{a_{\mathrm{eq}}}+1}\left(\frac{a}{a_{\mathrm{eq}}}-2\right)\right]\,,$ (A2)
where $a_{\mathrm{eq}}\equiv\Omega_{r}/\Omega_{\mathrm{M}}$ is the cosmic scale factor at matter-radiation equality.
Taking the ratio of $t$ at two different values of $a$ allows us to compute the ratio between the age of the universe at matter-radiation equality and recombination.
$\frac{t_{\mathrm{rec}}}{t_{\mathrm{eq}}}~{}=~{}\frac{2+\sqrt{\frac{a_{\mathrm{rec}}}{a_{\mathrm{eq}}}+1}\left(\frac{a_{\mathrm{rec}}}{a_{\mathrm{eq}}}-2\right)}{2-\sqrt{2}}\,,$ (A3)
where ‘eq’ subscripts refer to quantities at the time of matter-radiation equality and ‘rec’ subscripts refer to recombination. In $\Lambda$CDM, $a_{\mathrm{rec}}=1/1090$ and $a_{\mathrm{eq}}\approx 1/3410$, so $a_{\mathrm{rec}}/a_{\mathrm{eq}}\approx 3.13$ (table 2 of Planck Collaboration VI, 2020). This implies $t_{\mathrm{rec}}/t_{\mathrm{eq}}=7.3$.
It is possible to invert Equation A2 and obtain an expression for $a\left(t\right)$ by squaring both sides, which leads to a cubic expression for $a/a_{\mathrm{eq}}$. Substituting $u\equiv\left(a/a_{\mathrm{eq}}-1\right)/2$, we get an expression of the form
$4u^{3}-3u~{}=~{}f\left(t\right)\,,$ (A4)
for some function $f\left(t\right)$ that we do not repeat for clarity. This can be solved using a trigonometric triple angle formula by substituting $u=\cos\theta$, which yields
$\cos\left(3\theta\right)~{}=~{}f\left(t\right)\,.$ (A5)
Combining these results, we get our final expression for the expansion history in the era when matter and radiation are the only appreciable components in the Friedmann equation:
$\frac{a}{a_{\mathrm{eq}}}~{}=~{}1+2\cos\left[\frac{1}{3}\operatorname{acos}\left(\frac{1}{2}\left[\frac{3H_{0}t\sqrt{\Omega_{\mathrm{M}}}}{2\,a_{\mathrm{eq}}^{3/2}}-2\right]^{2}-1\right)\right]\,.$ (A6)
Our derivation remains valid if the argument of the central $\operatorname{acos}$ function becomes $>1$, but then we need to use the generalised definition of $\cos$ and $\operatorname{acos}$ for complex numbers. This is equivalent to setting $\cos\to\cosh$ and $\operatorname{acos}\to\operatorname{acosh}$ in Equation A6 and using only real quantities. We can alternatively use this equation as shown, in which case the result of the $\operatorname{acos}$ operation smoothly transitions from the real axis to the imaginary axis of the complex plane, but the $\cos$ operation then returns us back to the real axis.

### v2

@@
 Cosmology (343) – Hubble constant (758) – Matter density (1014) – Cosmological parameters from large-scale structure (340) – Stellar ages (1581)
-journal: The Open Journal of Astrophysics
+journal: Astronomy
 […]
@@
 We assume its fractional uncertainty is the same as that in $\Omega_{\mathrm{M}}h$.
+We use their study only to constrain this product rather than follow their approach of breaking the degeneracy using a prior on $\Omega_{\mathrm{M}}$, since this often comes from the same information already considered in Section 2.1, which we need to avoid double-counting.
 2.3 Stellar ages in the Galactic halo
@@
 This in turn would raise the expected cosmic variance in $H_{0}$ to perhaps 3 km/s/Mpc, which is sufficient to solve the Hubble tension as a $2\sigma$ local fluctuation.
-Alternatively, a modification to the expansion history at late times may have little impact on $t_{\mathrm{U}}$ and the angular diameter distance to the CMB.
-If the modification is carefully tuned, it could enhance the present expansion rate enough to solve the Hubble tension while not much affecting the other observables.
-Such a modification may arise from dark energy not being a pure cosmological constant (e.g., Harko, 2023; Yao & Meng, 2023).
+An interesting aspect of Figure 1 is that without the age constraint shown in blue, there is excellent agreement between the local $cz^{\prime}$ and the constraints from UCS and $k_{\mathrm{eq}}/h$.
+This suggests an early time resolution to the Hubble tension that allows the CMB anisotropies to be fit at higher $h$.
+The problem is the absolute ages of old stars (Cimatti & Moresco, 2023) and the Galactic disc as a whole (Xiang et al., 2024).
+It is also possible to constrain cosmology using relative stellar ages by treating stellar populations as cosmic chronometers (CCs; Moresco et al., 2018, 2020).
+The idea of the CC technique is that a ‘red and dead’ galaxy formed its stars in a rapid burst at very early times, with the stellar population passively evolving ever since.
+Since stars of different masses have different lifetimes, we may observe stellar populations in such galaxies at different redshifts to determine the relative age between those redshifts.
+This is similar to observing the main sequence turnoff in a globular cluster, but using relative rather than absolute ages.
+This can make the results less sensitive to evolutionary phases that might be hard to model, provided the model at least captures how the duration of such phases changes with stellar mass.
+The use of a much larger number of stars also reduces sensitivity to individual stars.
+The results obtained with CCs are almost entirely from $z>0.3$, making them largely insensitive to the very low redshifts at which a local void might play a role.
+This allows the CC method to constrain the background cosmology in a somewhat different way to nearby stars.
+The slope of the age-redshift relation can only constrain some combination of $H_{0}$ and $\Omega_{\mathrm{M}}$, so Cogato et al.
+(2024) combined CCs with uncalibrated SNe, BAOs, and $\gamma$-ray bursts (GRBs) treated as UCS, thereby obtaining $H_{0}=67.2^{+3.4}_{-3.2}$ km/s/Mpc.
+This is much more consistent with the Planck value than the local $cz^{\prime}$, which is $\approx 2\sigma$ higher.
+The good agreement between the constraints from CCs and from the absolute ages of ancient Galactic stars provides additional confidence that $H_{0}$ is below the local $cz^{\prime}$, at least if we assume an FLRW cosmology which is almost homogeneous today on the scales involved in the local $cz^{\prime}$ measurement.
+It is therefore difficult to solve the Hubble tension through modifications to the physics at very early times, even if a good fit to the CMB can be retained (a non-trivial task; see Vagnozzi, 2021).
+Thus, the most promising way to retain the usual assumption that $cz^{\prime}=H_{0}$ is to modify the expansion history at late times.
+This may have little impact on $t_{\mathrm{U}}$ and the angular diameter distance to the CMB.
+If the modification is carefully tuned, it could enhance the present expansion rate enough to solve the Hubble tension while not much affecting the other observables (Tiwari et al., 2024).
+Such a modification may arise from dark energy not being a pure cosmological constant (e.g., Harko, 2023; Yao & Meng, 2023; Rezazadeh et al., 2024).
 The main problem with this is that a larger $H_{0}$ requires a higher energy density according to the Friedmann equations, which would mean that the dark energy density rises with time (Dahmani et al., 2023).
@@
 where $a$ is the cosmic scale factor with value 1 today, an overdot denotes a time derivative, and $\Omega_{\mathrm{M}}$ ($\Omega_{\mathrm{r}}$) is the present fraction of the cosmic critical density in matter (radiation).
-We can solve this by separating the variables and using the substitution $a=a_{\mathrm{eq}}/\tan^{2}\theta$.
+We can solve this by separating the variables and using the substitution $a=a_{\mathrm{eq}}\sinh^{2}\theta$.
 Applying the usual boundary condition that $a=0$ when $t=0$ gives the following result:
@@
 This is equivalent to setting $\cos\to\cosh$ and $\operatorname{acos}\to\operatorname{acosh}$ in Equation A6 and using only real quantities.
-We can alternatively use this equation as shown, in which case the result of the $\operatorname{acos}$ operation smoothly transitions from the real axis to the imaginary axis of the complex plane, but the $\cos$ operation then returns us back to the real axis.
+If we instead use this equation as shown, the result of the $\operatorname{acos}$ operation smoothly transitions from the real axis to the imaginary axis of the complex plane, but the $\cos$ operation then returns us back to the real axis.